Mary is standing at the top of a hill admiring the view. The hill is at a 2.90° tilt. 29.0 m down the hill an old man lets go of a stroller with a baby inside it. The stroller starts to roll downhill with its frictionless wheels. Mary starts to run after it with her best acceleration of 2.90 m/s2. How far has the stroller rolled before she catches it.


1
Expert's answer
2021-09-18T14:59:04-0400

On the one hand, the distance Mary should run to catch the stroller:


"D=\frac{a_MT^2}2."

On the other hand, it is the distance the stroller covered plus x = 29 meters:


"D=d+x,\\\space\\
d=\frac{a_sT^2}2=\frac{gT^2\sin\theta}2 ,\\\space\\
D=\frac{gT^2\sin\theta}2 +x."

Equate D and find the time:


"\frac{a_MT^2}2=\frac{gT^2\sin\theta}2+x,\\\space\\
T^2=\frac{2x}{a_M-g\sin\theta}."

Therefore:


"D=\frac{xa_M}{a_M-g\sin\theta}=35\text{ m}."